Two small conducting and identical spheres A and B have charges -25 nC and +15 nC, respectively. They are separated by a distance of 0.02 m. (a) What is the magnitude of the electric force between the two spheres? Is this force attractive or repulsive? (b) The spheres are then allowed to touch each other and then separated. What is the magnitude of the force between the two spheres? Is this a repulsive force or an attractive force?



1
Expert's answer
2022-02-15T13:35:09-0500

Given:

"q_1=-\rm 25\:nC"

"q_2=\rm 15\:nC"

"r=\rm 0.02\:\rm m"


(a) The Coulomb law says



"F=k\frac{q_1q_2}{r^2}"

Hence



"F=9*10^9*\frac{25*10^{-9}*15*10^{-9}}{0.02^2}=8.4*10^{-3}\:\rm N"


(b) attraction

(c) the new charge of spheres

"q_1'=q_2'=\frac{q_1+q_2}{2}=-5\: \rm nC"

"F=9*10^9*\frac{5*10^{-9}*5*10^{-9}}{0.02^2}=5.6*10^{-4}\:\rm N"

Repulsive force.